OpenStudy (anonymous):

if 6a-8b=0 and c=12b, the ratio of a to c is equivalent to the ratio of one to what number?

OpenStudy (anonymous):

6a-8b = 0 6a = 8b 3a=4b a = 4/3b c = 12b a/c = (4/3)/12 = 1/9 a:c = 1:9

OpenStudy (anonymous):

solve for b b = c / 12 plug into 6a-8b=0 6a- (8/12) c = 0 6a = (2/3) c a/c = 1/9 ratio of one to 9